The problems associated with operating an effective "block list" system in an increasingly hostile environment

Robert Gallagher MSc in Security and Forensics School of Computing Dublin City University

February 10, 2004

Abstract

Bulk email, commonly referred to as ’Spam’, is a widely acknowledged problem. Several solutions have been proposed which range from placing levies on email in order to undermine the bulk emailing business model to including a new ’spam’ field in email message headers. These solutions would be very difficult to implement on top of the current email system, because of its widespread deployment. They also assume cooperation from the bulk emailers.

Anti-spam ’blocklist’ sites have taken a more direct approach by providing email users and Internet Service Providers with databases of Internet hosts known to harbour spammers. However, as bulk emailers have become more sophisticated in their techniques these block list sites are falling under attack.

This practicum will investigate the problems faced by blocklist systems and propose a solution based around a distributed blocklist system.